# ODMixer: Fine-grained Spatial-temporal MLP for Metro Origin-Destination Prediction

### v2

@@ IV-B2 Embedding Layer
$\boldsymbol{H_{i}}=\boldsymbol{W^{E}}\mathrm{OD}_{i}$ (4)
where $i=1,\cdots,N^{2}$, $\boldsymbol{W^{E}}\in\mathbb{R}^{d\times T}$ is learnable weights of Embedding layer, $\boldsymbol{H}\in\mathbb{R}^{N\times N\times d}$ represents the obtained feature representation after the Embedding Layer, $\mathrm{OD}$ denotes the data containing all OD pairs, and $\mathrm{EmbeddingLayer}$ is one-layer MLP. This approach differs from the previous methods that treat the OD matrix as a whole. Instead, we emphasize modeling each OD pair independently, to effectively capture the characteristics and dynamics of each OD pair.
@@
IV-C Origin-Destination Interact Module (ODIM)
ODIM conducts spatio-temporal interaction of features encoded by the Embedding Layer. It comprises two modules: Channel Mixer, which considers the short-term temporal relations of features within OD pairs, and Multi-view Mixer, which models the relations among OD pairs.
IV-C1 Channel Mixer
@@
IV-C2 Multi-view Mixer
@@
To address this issue, we introduce the Multi-view Mixer module by analyzing relations among OD pairs from multiple perspectives. Recognizing that each OD pair represents the flow between two stations, which is notably correlated with the origin and destination, we propose the Multi-view Mixer module, consisting of two distinct parts: Origin Mixer and Des Mixer.
@@
IV-D Bidirectional Trend Learner (BTL)
@@ IV-E1 Loss Function
$L_{prev}$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{d-1}-OD_{t+1}^{d-1}|}{N^{2}}$
$L_{cur}$ $=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{d}-OD_{t+1}^{d}|}{N^{2}}$
@@ IV-E2 Model Complexity
ODIM consists of two components: Channel Mixer and Multi-view Mixer. For Channel Mixer, the computational cost primarily arises from matrix multiplication in the linear layer, with intermediate dimensions being $2d$. This results in approximately $4N^{2}d^{2}$ operations. Because the $d$ is chosen independent of $N$, the time complexity of the Channel Mixer is $O(N^{2})$.
@@ V-A1 Dataset Description
We conduct experiments on two real-world datasets [5], as shown in Table II. HZMOD records transaction data for 80 stations from January 1, 2019, to January 25, 2019. SHMOD records transaction data for 288 stations from July 1, 2016, to September 30, 2016. These transaction data are segmented into 15-minute interval, and the complete dataset is partitioned into training, validation, and test set following a specific ratio.
@@
TABLE III: Performance of OD prediction on HZMOD dataset.
Models MAE ($\downarrow$) RMSE ($\downarrow$) wMAPE ($\downarrow$)
HA 1.355 2.917 48.354%
LSTM 1.387 3.458 49.500%
GRU 1.427 3.593 50.906%
Graph WaveNet 1.717 4.431 62.489%
DCRNN 1.269 2.944 46.203%
STG2Seq 1.578 4.355 56.302%
PVCGN 1.241 2.697 44.290%
DGSL 1.244 2.906 45.269%
Informer 1.272 2.756 45.393%
HIAM 1.196 2.581 42.690%
ODMixer (Ours) 1.131 2.367 40.358%
@@ V-A3 Quantitative Evaluation Metrics
In our study, we have selected three widely employed performance metrics within the domain of traffic flow prediction to serve as the benchmarks for evaluating the model’s effectiveness. These metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Weighted Mean Absolute Percentage Error (wMAPE).
@@ V-B Comparison with State-of-the-Art Methods
We conduct a comparative analysis between our proposed model and a selection of classical as well as contemporary models. STCNN [9] and ST-VGCN[10] that consider OD pairs are included because they have no publicly accessible codes and have not been evaluated on identical datasets.
@@
(a) Impact of Number of Layer
(b) Impact of Number of Dimension
Figure 4: Impact of Hyper-parameters. The two figure illustrate the impact of different parameters on prediction results across two datasets. Figure (a) displays the effect of varying the number of layers on model performance, while Figure (b) shows the influence of different feature dimensions on model performance.
We have conducted experiments on two datasets, and the experimental results for all methods are presented in Table III and IV. Firstly, the simplest method, HA, can partially reflect the flow range of the training set by directly averaging it. However, lacking generalization ability, it fails to capture temporal variations in traffic. Consequently, its performance tends to be inferior to that of ODMixer, with the performance gap widening as the dataset size increases. After analysis, it is evident that the performance of LSTM, GRU, Graph WaveNet, and STG2Seq models is inferior to that of HA, potentially due to their oversimplified design that neglects the spatial-temporal dependencies among metro stations. The incorporation of spatial-temporal relations in models like DCRNN and Informer has resulted in better performance. DGSL outperforms other baselines on the SHMOD dataset because it learns the graph structure rather than using the provided stricture, which allows for a more effective capture of relations on complex dataset. PVCGN, which considers three types of graphs and comprehensively accounts for the interstation relations, exhibits good performance across both datasets. HIAM, which integrates OD and DO information and considers the incompleteness of OD matrices, achieves commendable results but performs less effectively than DGSL on the SHMOD dataset, possibly because the large scale of SHMOD requires a more fine-grained approach than can be provided by neighborhood graph aggregation alone to capture the complex interactions among numerous stations.
@@
In contrast, ODMixer models the OD matrix with fine granularity from the perspective of OD pairs. Subsequently, it leverages the spatio-temporal relations among OD pairs through the ODIM module and integrates long-term temporal dynamics with the BTL module, thereby achieving superior performance.
@@ V-C Ablation Study
The results are presented in the Table V and VI. ODMixer-w/o CM exhibits inferior performance compared to ODMixer, signifying the effectiveness of Channel Mixer in modeling short-term temporal relations of OD pairs. ODMixer-w/o MM performs less effectively than both ODMixer-w/o OM and ODMixer-w/o DM, highlighting the necessity of both Origin and Destination perspectives. The results indicate that these two perspectives complement each other in considering the relations among OD pairs. ODMixer-w/o BTL and ODMixer-w/o PB show some performance losses, suggesting that accounting for the temporal similarity of metro flows can enhance the model’s performance. Due to the difficulty of accurately modeling long-term dependencies for large-scale metro data, ODMixer performs slightly worse than ODMixer-w/o BTL in RMSE metrics. These experimental findings affirm that all modules contribute to the ODMixer.
@@ V-E Efficiency Analysis
TABLE VIII: Efficiency study of ODMixer on SHMOD dataset
@@
In this subsection, we quantify the number of parameters for the various models. To ensure a fair comparison, all models are evaluated on a single NVIDIA GeForce RTX 2080 Ti GPU with a consistent batch size. The batch size is 8 for SHMOD and 32 for HZMOD.
Table VII and VIII present a comparative analysis of all methods. The LSTM, GRU, Graph WaveNet, and STG2Seq models exhibit poorer performance than HA, despite their shorter training and inference times. Despite Informer’s short training and inference times on the HZMOD dataset, it has a substantial number of parameters. Regarding the SHMOD dataset, Informer cannot be trained , even when the batch size is 1. Furthermore, the PVCGN model is unable to be trained on a single GPU with a batch size of 32 for HZMOD, necessitating a reduction to 16 for testing. Despite its good performance, the PVCGN model requires significantly longer training and inference times. In contrast, the HIAM model, which serves as the best-performance baseline, has a higher number of parameters, as well as longer training and inference times compared to ODMixer. Our proposed ODMixer not only achieves superior performance but also maintains reasonable training and inference times.
@@ V-F Validity Verification
TABLE IX: Performance of OD prediction on HZMOD dataset.
@@
ODMixer 40.609% 6.75 2.15 2 16 32 1
$\Delta$ 14,37% ×$\times 2.82$$\times 1.82$
@@
ODMixer 40.781% 8.04 2.59 3 16 32 1
$\Delta$ 13.24% ×$\times 3.46$$\times$ 1.81 - - - -
@@
The experiments show that the performance of ODMixer is improved by 14.37% and 13.24% at layers 2 and 3 compared to ODTrans, respectively. Additionally, by comparing training and testing times, we found that ODMixer’s training time and testing time at layer 2 were accelerated by 2.82 times and 1.82 times, respectively. At layer 3, training and testing times were accelerated by 3.46 times and 1.81 times, respectively. These results demonstrate that the ODMixer module, designed specifically for the metro OD prediction task, effectively captures the spatiotemporal relations in traffic. This significantly enhances the model’s operational efficiency, reduces the required space, and greatly improves its deployability.
@@ V-G Case Study
Figure 5: Case for SHMOD Dataset. It shows the prediction results of the DGSL and ODMixer models from September 10 to 13, with the 12th being a Monday. It can be observed that when traffic flow changes rapidly, the performance of DGSL is significantly worse than that of ODMixer.
@@
In this subsection, we use a practical example to demonstrate why our ODMixer achieves better results.
@@
As shown in Fig. 5, the OD traffic on Saturday and Sunday is minimal and stable, allowing both DGSL and ODMixer to fit the real traffic well. However, during the peak periods of Monday and Tuesday, the traffic between OD pairs changes rapidly. Although DGSL senses the growth and attenuation of traffic, its estimated peak value remains small. Conversely, ODMixer quickly detects these rapid changes and accurately estimates the peak value of the OD pair’s traffic. This is mainly due to ODMixer’s input modeling based on the OD pair perspective. The Channel Mixer captures the short-term temporal relations, the Multi-view Mixer learns the spatial relations between OD pairs, and BTL senses long-term traffic similarity. These designs enable ODMixer to handle rapid changes and complex relations in traffic effectively, significantly improving the model’s prediction accuracy and adaptability.

### v4

@@ IV-D Embedding Layer
${\color[rgb]{0,0,0}\boldsymbol{H}(i,j)=\boldsymbol{W^{E}}\mathrm{OD}_{t-T+1:t}(i,j)}$ (4)
where $i=1,\cdots,N,j=1,\cdots,N$, $\boldsymbol{W^{E}}\in\mathbb{R}^{d\times T}$ is learnable weights of Embedding layer, $\boldsymbol{H}\in\mathbb{R}^{N\times N\times d}$ represents the obtained feature representation after the Embedding Layer. This approach differs from the previous methods that treat the OD matrix as a whole. Instead, we emphasize modeling each OD pair independently, to effectively capture the characteristics and dynamics of each OD pair. After performing embedding layer on both Prev and Cur branches, we obtain the feature $\boldsymbol{H_{\mathrm{prev}}},\boldsymbol{H_{\mathrm{cur}}}\in\mathbb{R}^{N\times N\times d}$.
@@
IV-E Origin-Destination Interact Module (ODIM)
For the metro OD prediction task, the spatio-temporal relations plays a crucial role. First, passenger flow information in adjacent time periods is correlated, indicating that current passenger flow can be influenced by flows in preceding and succeeding periods. Second, each metro station has unique attributes, and stations with similar attributes typically exhibit similar passenger flow trends. Therefore, by learning and capturing the spatial relations of OD pairs, the model’s prediction performance can be significantly enhanced, leading to more accurate predictions of future passenger flow distribution. To address this, we designe the ODIM module to effectively capture the spatio-temporal relations of OD pairs. ODIM conducts spatio-temporal interaction of features encoded by the Embedding Layer. It comprises two modules: Channel Mixer, which considers the short-term temporal relations of features within OD pairs, and Multi-view Mixer, which models the relations among OD pairs.
IV-E1 Channel Mixer
@@
IV-E2 Multi-view Mixer
@@
To address this issue, we introduce the Multi-view Mixer module by analyzing relations among OD pairs from multiple perspectives. Recognizing that each OD pair represents the flow between two stations, which is correlated with the origin and destination, we propose the Multi-view Mixer module, consisting of two distinct parts: Origin Mixer and Des Mixer.
@@
IV-F Bidirectional Trend Learner (BTL)
@@ IV-G1 Loss Function
${\color[rgb]{0,0,0}L_{prev}}$ ${\color[rgb]{0,0,0}=}{\color[rgb]{0,0,0}\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{d-1}(i,j)-OD_{t+1}^{d-1}(i,j)|}{N^{2}}}$
${\color[rgb]{0,0,0}L_{cur}}$ ${\color[rgb]{0,0,0}=}{\color[rgb]{0,0,0}\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}|\widetilde{OD}_{t+1}^{d}(i,j)-OD_{t+1}^{d}(i,j)|}{N^{2}}}$
@@ IV-G2 Model Complexity
ODIM consists of two components: Channel Mixer and Multi-view Mixer. For Channel Mixer, the computational cost primarily arises from matrix multiplication in the linear layer, with intermediate dimensions being $2d$. This results in approximately $4N^{2}d^{2}$ operations. Since $d$ is chosen independent of $N$, the time complexity of the Channel Mixer is $O(N^{2})$.
@@ V-A1 Dataset Description
We conduct experiments on two real-world datasets [6], as shown in Table II. HZMOD records transaction data for 80 stations from January 1, 2019, to January 25, 2019. SHMOD records transaction data for 288 stations from July 1, 2016, to September 30, 2016. These transaction data are segmented into 15-minute interval, and the complete dataset is partitioned into training, validation, and test set following a specific ratio. Compared with previous subway datasets, these two datasets contain billions of transaction records from the Shanghai and Hangzhou subway systems in China, fully reflecting various real-world scenarios. The SHMOD dataset, in particular, includes 288 stations, resulting in 288*288 OD pairs, a substantial number. This scale introduces significant differences in the OD matrix, making it an excellent benchmark for assessing the model’s ability to handle large and complex OD matrices.
@@ V-A3 Quantitative Evaluation Metrics
We have selected three widely employed performance metrics within the domain of traffic flow prediction for evaluating the model’s effectiveness. These metrics include Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Weighted Mean Absolute Percentage Error (wMAPE). MAE is simple and intuitive, insensitive to outliers, suitable for scenarios that require robustness. RMSE penalizes large errors, provides information about the distribution of prediction errors, and is suitable for scenarios that need to emphasize the severity of errors. wMAPE expresses errors in percentage form, suitable for comparison of datasets of different sizes, and can be weighted to reflect the importance of data points.
@@ V-B Comparison with State-of-the-Art Methods
We conduct a comparative analysis between our proposed model and a selection of classical as well as contemporary models. STCNN [10] and ST-VGCN[11] that consider OD pairs aren’t included because they have no publicly accessible codes and have not been evaluated on identical datasets.
@@
  • HSTN[55]: HSTN designs a HSM to capture three types of spatial relations and a HTM to quantify the influence of the input sequence on the target result.
  • MVPF[5]: MVPF considers multiple views of real-time traffic information including individual station and cross-station flow information.
  • C-AHGCSP[56]: C-AHGCSP completes unfinished trips based on real-time mobility evolution and inter-station time costs to enhance dynamic adaptability and considers the destination distributions of the passengers departing from a station are correlated with other stations sharing similar attributes.
TABLE III: Performance of OD prediction on HZMOD dataset.
Models MAE ($\downarrow$) RMSE ($\downarrow$) wMAPE ($\downarrow$)
HA 1.355 2.917 48.354%
LSTM 1.387 3.458 49.500%
GRU 1.427 3.593 50.906%
Graph WaveNet 1.717 4.431 62.489%
HSTN 1.692 5.558 60.392%
MVPF 1.805 4.749 64.414%
DCRNN 1.269 2.944 46.203%
STG2Seq 1.578 4.355 56.302%
DGSL 1.244 2.906 45.269%
Informer 1.272 2.756 45.393%
PVCGN 1.241 2.697 44.290%
C-AHGCSP 1.239 2.694 44.200%
HIAM 1.196 2.581 42.690%
ODMixer (Ours) 1.131 2.367 40.358%
@@
(a) Impact of Number of Layer
(b) Impact of Number of Dimension
Figure 4: Impact of Hyper-parameters. The two figure illustrate the impact of different parameters on prediction results across two datasets. Figure (a) displays the effect of varying the number of layers on performance, while Figure (b) shows the influence of different feature dimensions on performance.
We have conducted experiments on two datasets, and the experimental results for all methods are presented in Table III and IV. We use the same data inputs as specified in the original papers when evaluating all baselines, provided the relevant data is available. For example, HIAM utilizes additional information, such as the Destination-Origin (DO) matrix, which we incorporate in our experiments. Firstly, the simplest method, HA, can partially reflect the flow range of the training set by directly averaging it. However, lacking generalization ability, it fails to capture temporal variations in traffic. Consequently, its performance tends to be inferior to that of ODMixer, with the performance gap widening as the dataset size increases. After analysis, it is evident that the performance of LSTM, GRU, Graph WaveNet, STG2Seq, HSTM and MVPF models is inferior to that of HA, potentially due to their oversimplified design that neglects the spatial-temporal dependencies among metro stations. C-AHGCSP addresses the issue of incomplete OD matrices by introducing a dedicated module for completion. However, its approach involves sequentially filling in each time interval step by step. And it primarily utilizes the flow information from the most recent moment and uses simple module to learn dependencies. Therefore it performs well in smaller-scale metro networks, it faces limitations when dealing with larger-scale metro networks The incorporation of spatial-temporal relations in models like DCRNN and Informer has resulted in better performance. DGSL outperforms other baselines on the SHMOD dataset because it learns the graph structure rather than using the provided structure, which allows for a more effective capture of relations on complex dataset. PVCGN, which considers three types of graphs and comprehensively accounts for the inter-station relations, exhibits good performance across both datasets. HIAM, which integrates OD and DO information and considers the incompleteness of OD matrices, achieves commendable results but performs less effectively than DGSL on the SHMOD dataset, possibly because the large scale of SHMOD requires a more fine-grained approach than can be provided by neighborhood graph aggregation alone to capture the complex interactions among numerous stations.
@@
In contrast, ODMixer models the OD matrix with fine granularity from the perspective of OD pairs. Subsequently, it leverages spatial-temporal relations among OD pairs through the ODIM module and integrates long-term temporal dynamics with the BTL module, thereby achieving superior performance.
@@ V-C Ablation Study
  • ODMixer-w/o OMP: Remove OD Matrix Processing module, eliminating the impact of unfinished orders.
@@
The results are presented in the Table V and VI. ODMixer-w/o OMP removes the OD Matrix Processing module, the Cur Branch receives only the incomplete OD matrix as input, without any completion processing. This incompleteness results in the model lacking critical information needed to capture the traffic characteristics at the current time, leading to an inaccurate reflection of the actual traffic situation. Consequently, the model struggles to make effective predictions, and its performance drops significantly. ODMixer-w/o CM exhibits inferior performance compared to ODMixer, signifying the effectiveness of Channel Mixer in modeling short-term temporal relations of OD pairs. ODMixer-w/o MM performs less effectively than both ODMixer-w/o OM and ODMixer-w/o DM, highlighting the necessity of both Origin and Destination perspectives. The results indicate that these two perspectives complement each other in considering the relations among OD pairs. The Multi-view Mixer’s effectiveness can be attributed to the fact that each OD pair is associated with both an origin and a destination, and the attributes of these locations influence the passenger flow dynamics. For example: If the origin is a residential area and the destination is a business district, there is typically a morning peak in passenger flow, with lower flows during other times. These patterns validate the effectiveness and rationale of modeling spatial relations from both dimensions. ODMixer-w/o BTL and ODMixer-w/o PB show some performance losses, suggesting that accounting for the temporal similarity of metro flows can enhance the model’s performance. Due to the difficulty of accurately modeling long-term dependencies for large-scale metro data, ODMixer performs slightly worse than ODMixer-w/o BTL in RMSE metrics. These experimental findings affirm that all modules contribute to the ODMixer.
V-D Hyper-parameters Analysis
@@ V-E Efficiency Analysis
TABLE VIII: Efficiency study of ODMixer on SHMOD dataset.
@@
In this subsection, we quantify the number of parameters for the various models. To ensure a fair comparison, all models are evaluated on a single NVIDIA GeForce RTX 2080 Ti GPU with a consistent batch size. The batch size is 8 for SHMOD and 32 for HZMOD, respectively.
Table VII and VIII present a comparative analysis of all methods. The LSTM, GRU, Graph WaveNet, STG2Seq, HSTN and MVPF models exhibit poorer performance than HA, despite their shorter training and inference times. Because C-AHGCSP completes the OD matrix step by step introduces additional computational overhead, it needs longer inference times. Despite Informer’s short training and inference times on the HZMOD dataset, it has a substantial number of parameters. Regarding the SHMOD dataset, Informer cannot be trained , even when the batch size is 1. Furthermore, the PVCGN model is unable to be trained on a single GPU with a batch size of 32 for HZMOD, necessitating a reduction to 16 for testing. Despite its good performance, the PVCGN model requires significantly longer training and inference times. In contrast, the HIAM model, which serves as the best-performance baseline, has a higher number of parameters, as well as longer training and inference times compared to ODMixer. Our proposed ODMixer not only achieves superior performance but also maintains reasonable training and inference times.
@@ V-F Validity Verification
TABLE IX: Validity Verification on HZMOD Dataset.
@@
ODMixer 40.609% 2.85 0.88 2 16 32 1
$\Delta$ 14,37% ×$\times 6.67$$\times 4.45$
@@
ODMixer 40.781% 3.75 0.98 3 16 32 1
$\Delta$ 13.24% ×$\times 7.41$$\times 4.80$
@@
(a) Case for SHMOD Dataset
(b) Case for HZMOD Dataset
Figure 5: Case for SHMOD and HZMOD Dataset. (a) shows the prediction results of the DGSL and ODMixer models from September 10 to 13, with the 12th being a Monday. (b) shows the prediction results of the HIAM and ODMixer models from January 21 to 24, with the 21th being a Monday. It can be observed that when traffic flow changes rapidly, the performance of HIAM and DGSL is significantly worse than that of ODMixer.
@@
The experiments show that the performance of ODMixer is improved by 14.37% and 13.24% at layers 2 and 3 compared to ODTrans, respectively. Additionally, by comparing training and testing times, we found that ODMixer’s training time and testing time at layer 2 were accelerated by 6.67 times and 4.45 times, respectively. At layer 3, training and testing times were accelerated by 7.41 times and 4.80 times, respectively. These results demonstrate that the ODMixer module, designed specifically for the metro OD prediction task, effectively captures the spatiotemporal relations in traffic. This significantly enhances the model’s operational efficiency, reduces the required space, and greatly improves its deployability.
@@ V-G Case Study
In this subsection, we use two practical examples to demonstrate why our ODMixer achieves better results.
@@
As shown in Fig. 5 (a), the OD traffic on Saturday and Sunday (Sep 10-11) is minimal and stable, allowing both DGSL and ODMixer to fit the real traffic well. However, during the peak periods of Monday and Tuesday (Sep 12-13), the traffic between OD pairs changes rapidly. Although DGSL senses the growth and attenuation of traffic, its estimated peak value remains small. Conversely, ODMixer quickly detects these rapid changes and accurately estimates the peak value of the OD pair’s traffic. This is mainly due to ODMixer’s input modeling based on the OD pair perspective. The Channel Mixer captures the short-term temporal relations, the Multi-view Mixer learns the spatial relations between OD pairs, and BTL senses long-term traffic similarity. These designs enable ODMixer to handle rapid changes and complex relations in traffic effectively, significantly improving the model’s prediction accuracy and adaptability. A similar situation is observed with HIAM and our ODMixer in Fig. 5 (b).
V-H Scalability
(a) Time curves of different L
(b) Time curves of different d
Figure 6: Comparasions between running time and various L, d.
Scalability tests are performed to assess the impact of various hyperparameters on model efficiency, with a primary focus on the number of layers $L$ and the hidden layer dimension $d$. First, we evaluate the effect of the number of layers. As shown in Fig. 6(a), both training and inference times increase as the number of layers grows, although the increase in inference time is relatively smaller. Considering the trade-off between performance (Fig. 4(a)) and computational cost, we select 5 layers to strike a balance between effectiveness and efficiency.
Next, we examine the effect of the hidden layer dimension, as depicted in Fig. 6(b). The results indicate that training time varies with the hidden layer dimension, while inference time remains relatively stable. Based on overall performance metrics (Fig. 4(b)), we choose a hidden layer dimension of 16. Overall, the model’s inference time exhibits minimal variation, and training time remains within acceptable limits. These findings confirm that ODMixer can be efficiently trained on larger datasets and deployed in real-time applications. Thus, ODMixer demonstrates strong stability and adaptability under different hyperparameter settings, making it highly suitable for real-time scenarios.
V-I Robustness
TABLE X: Prediction error of Models on original data(1st row), Gaussian-noise polluted data(2nd row) and the relative increment ratio of the error(3rd row).
Model HZMOD SHMOD
MAE RMSE wMAPE MAE RMSE wMAPE
HIAM 1.196 2.581 42.690% 0.441 1.226 60.268%
+ $\mathcal{N}(0,1)$ 1.214 2.638 43.337% 0.454 1.279 62.049%
+$\Delta$errors 1.51% 2.21% 1.52% 2.95% 4.32% 2.96%
ODMixer 1.131 2.367 40.358% 0.408 1.096 55.691%
+ $\mathcal{N}(0,1)$ 1.136 2.380 40.538% 0.413 1.122 56.384%
+$\Delta$errors 0.44% 0.55% 0.45% 1.23% 2.37% 1.24%
In real-world metro systems, hardware or software failures such as sensor malfunctions or network issues may result in missing order data, leading to discrepancies between the obtained traffic data and the actual data. To assess the robustness of the model under such conditions, we evaluate its performance under noisy and missing data scenarios. For the noisy data case, we add Gaussian noises into the raw data of test dataset and test the model on the noisy test dataset using the trained model[57]. The results in Table X show the increasing errors of ODMixer are much less than SOTA for the noise data, verifying the robustness of ODMixer.
For the missing data scenario, we perform spatiotemporal masking on the test dataset, randomly masking a portion of the data to simulate potential missing data situations. The results in Fig.7 show the prediction errors is gradually increasing when mask ratio is increasing. When the mask ratio is relatively low (not greater than $30\%$), ODMixer still maintains good performance compared to the SOTA models, demonstrating that ODMixer has some resilience to missing data. Through experiments in both noisy and missing data scenarios, we further demonstrate that ODMixer exhibits robust performance, effectively handling common issues of noise and data missing in real-world applications. This ensures the model’s stability and reliablity in real-world metro environments. Traffic prediction robustness focuses on enhancing a model’s resilience to noise and handling missing data, which often referred to as the imputation task. Several studies have explored model robustness [57, 58]. Future work could integrate noise resilience and imputation modules to preprocess noisy or incomplete data, improving prediction quality.
(a) HZMOD
(b) SHMOD
Figure 7: MAE curve for different missing ratio on HZMOD and SHMOD.
V-J Cross-City Generalizability
We conduct a comprehensive experiment using two real-world datasets with distinct characteristics. The results demonstrate that our model can effectively capture the complex spatiotemporal relations inherent in different datasets and is well-suited for deployment in metro systems with varying characteristics. To further validate the generalization capability of our model, we design the cross-city experiment, where the model is trained on the metro network of one city and fine-tuned and tested on a different city’s network with distinct characteristics. However, due to differences in the number of stations across cities, the number of OD pairs also varies.
Therefore, to ensure the model operates properly, we map the number of stations at the input stage accordingly. We adopted two efficient mapping methods, OAda and TAda, to address differences in the number of stations between cities. These methods use Adapters to map the number of stations from the source city to the target city ($N_{source\_city}\rightarrow N_{target\_city}$), where $N_{source\_city}$ and $N_{target\_city}$ represent the numbers of metro stations in the source and target cities, respectively (e.g., HZMOD has 80 stations, while SHMOD has 288). Each Adapter is implemented as a two-layer MLP. In the OAda method, the origin and destination share a single Adapter, whereas in the TAda method, the origin and destination each use separate Adapters. We test the performance of two fine-tuning approaches: full model fine-tuning and Adapter-layer-only fine-tuning. The experimental results, shown in the Table XI, indicate that full model fine-tuning outperforms Adapter-layer-only fine-tuning and achieves better results than the current SOTA methods. Interestingly, Adapter-layer-only fine-tuning performs comparably to SOTA and even slightly surpasses it under the TAda configuration. Comparing the two mapping methods, TAda generally outperforms OAda, especially when only the Adapter layer is fine-tuned. These results demonstrate that ODMixer has strong cross-city generalization and adaptability, effectively transferring knowledge between metro networks with different attributes.
Recent research has introduced methods specifically designed to enhance the generalization of traffic prediction models [59, 60]. These approaches aim to improve a model’s ability to transfer knowledge across datasets with varying attributes. The primary objective is to train models on large, data-rich datasets and fine-tune them on smaller, domain-specific datasets, addressing challenges related to data scarcity. In future work, ODMixer could be extended to develop a more generalized metro OD prediction model.
TABLE XI: Cross-City performance of different methods. FF: Full Fine-tuning. OFA: Only Fine-tuning Adapter. OAda: One Adapter. TAda: Two Adapters.
Method SHMOD $\rightarrow$ HZMOD HZMOD $\rightarrow$ SHMOD
MAE RMSE wMAPE MAE RMSE wMAPE
FF OAda 1.169 2.481 41.700% 0.436 1.163 59.553%
TAda 1.169 2.462 41.732% 0.431 1.148 58.929%
OFA OAda 1.212 2.580 43.243% 0.449 1.231 61.359%
TAda 1.195 2.532 42.632% 0.435 1.165 59.357%